OpenStudy (anonymous):

y=.0375cos(.45x+5.4t) how far does the crest of the wave travel in 1.16s?

OpenStudy (anonymous):

Equation of a wave y=Asin(kx+wt) here wave velocity=w/k hence v=5.4/4.5 distance=speed * time=(54/45)*16